OpenStudy (anonymous):

whats the anti derivative of (2/x)^1/2? I'm a bit puzzled

OpenStudy (turingtest):

converting to fractional exponents makes the problem easy in cases like this\[\int(\frac2x)^{1/2}dx=\sqrt2\int x^{-1/2}dx\]take it from here